OpenStudy (anonymous):

Parallel to 3x – y = -5 and passing though (-1,0)

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

m=-3 y-0 = -3(x-1) 3x + y = 1

OpenStudy (anonymous):

no sorry thats wrong

OpenStudy (anonymous):

the slope of the first is 3

OpenStudy (anonymous):

y-y1 = m(x-x1) y=3(x+1) y = 3x + 3

OpenStudy (anonymous):

thanks